10. Graph by picking a few polar points. a) r = 2 - 2 cose and b) graph r = .3sine (calculator is for calculation only)
Answer #1

Ye 2 -2_CoS 9 r 0° a-acos(G) = 2-341) = 0 90 a-co%(96) - -300)= 2 186 2-3Cor(185) = 4-2-1) =y = 8-2 (0) = 4 2- aco (ฉาย) 96 (48 0 188 0 0 어 - 2 3 a78